For a Saturday matinee, adult tickets cost $3.50 and kids under 12 pay only $3.00. If 77 tickets are sold for a total of $254, how many of the tickets were adult tickets and how many were sold to kids under 12?
Answer by ewatrrr(24785)

Hi
$3.50x + $3.00(77-x) = $254 (x the number of adult tickets)
x = (254 - 231)/.50 = 46 Adult, kids 31 77-46
check answer
$3.50*46 + $3.00*31 = ____